OpenStudy (anonymous):

Prove that the sequence \((a_n)_{n\in \mathbb{n}}\) is bounded iff the sequence \((|a_n|)_{n\in \mathbb{n}}\) is bounded from above.

OpenStudy (freckles):

for one way... you can suppose |a|<=p for all natural n and p is greater than or equal to 0 |a|<=p implies something about the sequence a

OpenStudy (freckles):

for the other way... I would use the fact that |x|=sqrt(x^2)

OpenStudy (anonymous):

"\(\Rightarrow\)", Let \(M_1\) be an upper bound for \((a_n)_{n\in \mathbb{N}}\), and \(M_2\) be an lower bound for \((a_n)_{n\in \mathbb{N}}\). Then, we have \(M_2\le a_n \le M_1\) for all n \(\in \mathbb{N}\). Take \(M = \max(|M_1|, |M_2|)\), then for all n \(\in \mathbb{N}\), we have, \(-M\le M_2\le a_n \le M_1 \le M\), i.e. \(|a_n| \le M\) Hence, \((|a_n|)_{n\in \mathbb{N}}\) is bounded from above if \((a_n)_{n\in \mathbb{N}}\) is bounded "\(\Leftarrow\)", Let \(M\) be an upper bound for \((|a_n|)_{n\in \mathbb{N}}\). Then, for all n \(\in \mathbb{N}\), we have \(|a_n|\le M\), that is \(-M\le a_n\le M\). Hence, \((a_n)_{n\in \mathbb{N}}\) is bounded if \((a_n)_{n\in \mathbb{N}}\) is bounded from above. Therefore, the sequence \((a_n)_{n\in \mathbb{N}}\) is bounded iff the sequence (\(|a_n|)_{n\in \mathbb{N}}\) is bounded from above. Does this proof look okay?

OpenStudy (freckles):

looks fine to me it actually looks more formal than what I have written

OpenStudy (anonymous):

We need a formal proof. :)

OpenStudy (zzr0ck3r):

looks great.